# Career Modeling with Missing Data and Traces

### v2

@@ 3.3.4 Substantive Results
Dependency with age and cohort. Figure 2 summarises information related to the probability of transitioning between states depending on age for model 3. It was computed by taking the inverse logistic transform of the sum of the intercept with the age spline coefficients. We first observe, that, around 0 years after admission, the probability of leaving for the private sector $\gamma^{0}$ is very low; this was to be expected, since individuals who never leave any trace (e.g. because they immediately resigned from ENA) are not included in the dataset. The probability $\gamma^{1}$ that an individual moves back from the private to the public sector is very high, but note that this concerns a very slim proportion of the population (less than .$7\%.$ up to three years). The median probability $\gamma^{0}$ then increases to $0.013$ where it remains more or less constant up to 25 years after admission. A stabilization occurs for $\gamma^{1}$ at around double the rate, at $0.025$. This is interesting as we observe there isn’t an important variation in the probability of leaving across the career, and the hypothesis of an overall geometric rate for leaving and coming back is quite realistic. We interpret the change of behavior 30 years after admission as artefactual, as almost no one in the population has had the time to arrive at this stage, which can also be seen with the widening of uncertainty.
@@
We now turn to our main quantities of interest, cohort-related coefficients, which are presented in the left pane of Figure 3. The view is quite contrasted between $\gamma^{0}$ and $\gamma^{1}$. On the one hand, it isn’t possible to conclude that the chance of leaving for the private sector $\gamma^{0}$ has significantly changed across the period, as the credible interval well includes zero ($\text{med}=-0.226,I_{90}=[-0.625,0.176]$). However, it is at least possible to state strongly that this coefficient has not increased across time, and has quite possibly decreased. On the other hand, we can state with fair confidence that $\gamma^{1}$ has increased over time, with $\text{med}=1.16$, ($I_{90}=[0.301,1.98]$). These results are robust to misspecification, as shown by checks with modular inference described in Appendix C. We thus conclude that on the one hand, the probability of leaving the public sector has not increased in the period, and may even have decreased; and that the probability of coming back has increased in the period, contrary to what we expected before conducting this study.
@@
Dependency with the electoral process. The posterior for the election time effects are reported in the right pane of Figure 3. We observe that the overall size of potential effects is far smaller than for period effects, and so should be interpreted cautiously. On the one hand, we conclude that elections have no impact on $\gamma^{0}$, as the credible interval is very well centered on zero ($\text{med}=-0.0212,I_{90}=[-0.249,0.201]$). On the other hand, however, despite important uncertainty, there seems to be a small positive effect of election time on $\gamma^{1}$ ($\text{med}=0.302,I_{90}=[0.000,0.302]$). In summary, electoral shifts probably do not influence the chance of high civil servants to leave for the private sector, but it might lead to higher chances of coming back for those who left.
3.3.5 Discussion of the substantive results
We conclude that exit rates from the state have not increased on the period for former ENA students, contrary to what we expected before conducting this study. If any change was to be inferred from data, it would rather be that elite civil servants have now a lower probability of leaving for the private sector than before, but it is possible to remain undecisive. Furthermore, the rate of former civil servants coming back to work in the public sector after leaving has significantly increased over the period. In other words, when civil servants leave for the private sector, they now have a higher chance of coming back. This is notably a contribution with respect to (Bouzidi et al., 2010), who stressed that they could not study such evolutions given the constraints of their data despite the demands of their journal referees, and showcases the advantages of studying larger groups via data augmentation.
In addition, if there seems to be an interaction between political majority changes with election years and revolving doors for this population, it seems to affect the chance of coming back from the private sector than the probability of leaving.
3.4 Study 2. Public-private paths across the state
Many social scientists have early insisted on the usefulness of adopting a comparative point of view when describing civil servants populations and public-private paths (Charle, 1987). We are able with our method to describe the heterogeneity of public-private paths across the State and how they interact with gender.
3.4.1 Population definition
Top administrators are divided between groups and services that are sometimes tied to certain special statutes. While recent regulatory changes have attempted to increase mobility between them and foster a common organizational culture, they remain useful categories to base our approach upon.
Our study focuses on the following groups: three inspection groups whose mission is to audit public organizations organized thematically around financial affairs (igf, Inspection Générale des Finances), social affairs (igas, Inspection Générale des Affaires Sociales), and administration (iga, Inspection Générale de l’Administration); the Cour des Comptes (ccomptes), the supreme financial jurisdiction with a similar financial auditing function; the Conseil d’État (ce), whose members serve both as judges in the Supreme Administrative Court and as legal advisors to the government; the préfets (cprefet), state administration representatives in regions and departments; ambassadors (cdiplo, Corps Diplomatique); administrators of the national statistical and economics office (insee) who can serve key roles in the definition of economic policies; and finally, directors and assistant directors of two powerful services of the Ministry of the Economy, treasury (dgtresor), and public finances (dgfip).
This selection is voluntarily diverse and encompassing, as one modeling goal is to describe how public-private mobility is a heterogeneous phenomenon. We describe in Appendix A the method we use to link individuals to each group. The total number of individuals in each group alongside the proportion of them matched to a digital profile is given in Appendix B. Note that some individuals are affiliated with multiple services because their affiliation may change over time, a technicality we discuss below. The number of missing values is greater than in study 1, reaching 75%, which makes data augmentation even more valuable.
3.4.2 Relevant applied literature and hypotheses
Different researchers have studied the trajectories of high civil servants in France by defining populations based on organizational belonging at a certain time point, especially by focusing on their past trajectory.
Rouban (2010) studied the finance general inspection (igf) on the 1958-2008 period. igf appears to have an important place in the revolving doors phenomenon, as they identified that over the 578 inspectors present in the organization between 1958-2009, a majority had wandered at least once in the private sector (62%). They identify a discontinuity around 1990, with a massification of individuals having done a business school before attending ENA and then igf. They also observed after 2000 a growing proportion of igf members who have had an experience in the private sector before coming back, which seems quite aligned with results obtained in our first study with a positive cohort effect for $\gamma^{1}$. However, we underlie that this study too is subject to bias because of the lack of accounting of right-censoring.
Kolopp (2021) conducted a study on the general direction of the treasury (dgtresor) in the 1965-2010 period. They exhibited a gendered behavior: men are more likely than women to leave for the most prestigious private industries (especially banking); on the other hand, they do not seem to show gender differences in the overall probability of leaving the public sector. They describe that differentiated practices generate unequal outcomes at different points in careers, for example with the stress on international mobility for career advancement and differences in professional networks. If we have reasons to assume these factors also play a significant role for other organizations, their strength may vary, and other factors might come into play. Even if we don’t differentiate within private sector activities, we will compare behavior between men and women across the different organizations of our sample.
@@ 3.4.2 Relevant applied literature and hypotheses
We attempt to build upon their results and propose a complete and systematic comparison across the different organizations we mentioned. This has value in the sense that it can qualify statements regarding the most prestigious organizations that are represented in these studies: are they really that special compared to other powerful services? For example, are the Finance Inspection and Treasury that much more affected by revolving doors compared to other services in the economy ministry or other inspections? We can also test statements linked to individual properties: do men have higher chances to leave for the private sector, and then come back?
@@ 3.4.3 Models
We encode organizational membership using a set of $D$ dichotomous variables $M_{t,1}^{i},...,M_{t,D}^{i}$ corresponding to our $D=11$ selected organizations, that we will include in the covariate matrices $W_{t}^{i}$. Contrary to the first study, we need to account for the fact that individuals can be affiliated to multiple organizations during their professional career. To remain consistent with our choice to focus on the main organizational identity of individuals, we impose that someone, at any given time, may only be linked to one organization in our set of $D$ organizations. This means we impose $\sum_{d=1}^{D}M_{t,d}^{i}=1$. If an individual is observed to become part of organization $d_{1}$ at time $t_{1}$ and then of organization $d_{2}$ at time $t_{2}$, we will have $M_{t,{d_{1}}}=1$ for $t_{1}\leq t<t_{2}$, and $M_{t,d_{2}}=1$ for $t\geq t_{2}$. This is a simplification, which has the advantage of enabling an analysis that is closest to what we could obtain by performing completely separate analyses for the different public organizations, even if it could be altered to study how having been the member of a given organization can affect behavior once being a member of another one included in the study.
We implement two configurations. Model 1 is a Markov switching model for which we only include an intercept for each group (no general intercept, and no reference organization), in addition to the autoregressive covariates for traces. The autoregressive covariates are interacted with the dummy of every organization. This is our reference model. Model 2 adds to the reference model a dichotomous variable encoding gender ($G^{i}=1$ if $i$ is a woman), for both $\gamma^{0},\gamma^{1}$ and $\lambda$.
@@ 3.4.4 Results
Reconstruction loss. The reconstruction loss obtained by the two models in this study is of $0.279$ and $0.265$, which is higher than in study 1. This can be explained because the inference task is more complex given the higher heterogeneity of the population, and is in line with what could be expected given the nature of the application.
@@
Comparison across organizations. We report the $5,50,$ and $95$ quantiles of the posterior for each of these models. Median estimates of coefficients for Model 1 are reported in Figure 4, alongside their 90% credible interval, expressed in raw probability. Bear in mind that these are average coefficients that do not correspond to local probabilities individuals may have to leave at any given point in time, which are modulated by the observed traces and the local $\lambda$’s. As coefficients for dgfip stand out from the rest in both estimate and uncertainty, we do not report them in the Figure but do so in the text below.
@@
Figure 4: Median and 90% credible interval of coefficients $\gamma^{0}_{\text{group}}$, $\gamma^{1}_{\text{group}}$ for individuals belonging to groups defined in subsection 3.4.1. Coefficients may be interpreted as average exit ($\gamma^{0}$) or come-back ($\gamma^{1}$) rates, yet it does not mean the model predicts transitions to happen at a constant rate.
@@
We can firstly observe and rank organizations depending on the chance individuals have to leave for the private sector when they are part of them. On the side of organizations with the highest departure rate, we find the General Inspection of Finances ($\gamma^{0}_{\text{igf}}=0.031$) and Prefects group ($\gamma^{0}_{\text{cprefet}}=0.029$). In an intermediary situation, we find the Diplomats group ($\gamma^{0}_{\text{cdiplo}}=0.022$), and the General Inspection of Administration ($\gamma^{0}_{\text{iga}}=0.020$) At the lower hand of the spectrum, we find the statistical office ($\gamma^{0}_{\text{insee}}=0.016$), the Cour des Comptes ($\gamma^{0}_{\text{ccomptes}}=0.015$), Treasury ($\gamma^{0}_{\text{dgtresor}}=0.014$), the General Inspection for Social Affairs ($\gamma^{0}_{\text{igas}}=0.013$), and the Conseil d’État ($\gamma^{0}_{\text{ce}}=0.011$). Not shown in the table, the General Direction for Public Finances has the lowest probability ($\gamma^{0}_{\text{dgfip}}=0.006,I_{90}=[0.004,0.009]$). This last very low coefficient means that there are almost no departures from this direction, and when that is the case, civil servants move first to another service included in this study. The overall ranking should be intuitive to the reader accustomed to the question – the model rightly identifies that individuals in the Treasury have a higher chance of leaving for the private sector than individuals in the Inspection for Social Affairs. A seemingly counterintuitive result is the position of igf and ccomptes: these two groups have a similar function (financial auditing of the state), but have very different positions in Figure 4; this observation is in fact aligned with an effect identified in the survey from Charle (1987).
@@
We observe interesting contrasts in the chance of individuals to come back working in public administration after being in the private sector. Inspection of Finances again clearly stand out, with ($\gamma^{1}_{\text{igf}}=0.024$). With varying levels of uncertainty, Conseil d’État ($\gamma^{1}_{\text{ce}}=0.020$), Inspection of social affairs ($\gamma^{1}_{\text{igas}}=0.019$), Treasury ($\gamma^{1}_{\text{dgtresor}}=0.018$), and Cour des comptes ($\gamma^{1}_{\text{dgtresor}}=0.016$), they appear in an intermediate situation. At the lower end of the spectrum, we observe the statistical office ($\gamma^{1}_{\text{insee}}=0.006$), inspection of administration ($\gamma^{1}_{\text{iga}}=0.004$), prefects ($\gamma^{1}_{\text{cprefets}}=0.003$), and diplomats ($\gamma^{1}_{\text{insee}}=0.002$).
@@
We thus conclude that the two rankings are not convergent and would even be negatively correlated if we excluded igf. However, we find it more accurate to describe this as forming three groups. First, igf constitutes a group on its own, standing out from the rest as having both high probabilities for leaving and coming back. Classical results claiming that moving into the private sector is part of the career advancement process hold well, and we further prove that this is quite distinctive to this organization. We identify a second group of organizations where the chance of leaving is relatively high, but the chance of coming back is rather low: this concerns the statistical office, the inspection of administration, diplomats and préfets. The situation is unsurprising for the first two, which are known to be closer to the private sector than other organizations. For préfets and diplomats, this might be because we study individuals who are simultaneously at the top of certain services but without as high position security as in other organizations (such as Conseil d’État, where people can keep their status): people (involuntarily) leaving could find themselves unable to come back. Finally, we find a group of organizations where the chance to leave is comparatively low and the chance to come back high: Conseil d’État, Inspection of Social Affairs, Cour des Comptes. One could also place Treasury in this group, but parameter uncertainty (especially on $\gamma^{1}$) requires caution. This concerns organizations for which members acquire life-long status that can help them come back after leaving; and also for which clear career advancement schemes exist in the public sector.
Interaction with gender
@@
The posterior of the odds ratio for women versus men can be found in Figure 6 as a way to summarise interest parameters for model 2. A distribution on the right means we have grounds to think women have a higher chance of leaving for the private sector than men, and vice-versa. A distribution centered on 1 indicates we should remain indecisive on the matter.
@@
Figure 5: Posterior distribution of $\frac{\gamma^{0}_{\text{group},W}/(1-\gamma^{0}_{\text{group},W})}{\gamma^{0}_{\text{group},M}/(1-\gamma^{0}_{\text{group},M})}$, the odds-ratio of the differential of exit rate between women and men expressed, for each of the nine groups in Study 2; the acronyms of the groups are defined in subsection 3.4.1. When the posterior is concentrated on the left-hand side, we can conclude that women in that group have a lower probability than men of exiting the public sector.
@@
Not all organizations have the same profile of differentiated practices between men and women, but a broad conclusion can nonetheless be inferred: women have less chance than men to leave for the private sector. For certain organizations, the difference is strong and identifiable: Cour des Comptes, Inspection of Administration, Inspection of Social Affairs (odds ratios below 0.7). The difference is smaller but still observable for Inspection of Finances, Préfets, and Diplomats. The situation is undecided for Conseil État and the Statistical Office, the latter suggesting a small positive effect on the side of women. Treasury here stands out in terms of uncertainty (as there are less than 200 individuals to work on and a small proportion of women), but the side of the effect would be rather positive for women.
3.5 Discussion on the substantive results
We find three profiles of organizations depending on the chance of leaving and the chance of coming back: one with high chances of leaving but low chances of coming back; one with low chances of leaving but high chances of coming back; and Inspection of Finances standing out from the rest with high chances of both to leave and to come back. Women have overall less chance to leave for the private sector than men, but the situation is not constant across public organizations.
4 General discussion

### v3

@@ 5.2.3 Substantive Results
Dependency with age and cohort. Figure 3 summarises information related to the probability of transitioning between states depending on age for model 3. It was computed by taking the inverse logistic transform of the sum of the intercept with the age spline coefficients. We first observe, that, around 0 years after admission, the probability of leaving for the private sector $\gamma^{0}$ is very low; this was to be expected, since individuals who never leave any trace (e.g. because they immediately resigned from ENA) are not included in the dataset. The probability $\gamma^{1}$ that an individual moves back from the private to the public sector is very high, but note that this concerns a very slim proportion of the population (less than $0.7\%$ up to three years). The median probability $\gamma^{0}$ then increases to $0.013$ where it remains more or less constant up to 25 years after admission. A stabilization occurs for $\gamma^{1}$ at around double that value, at $0.025$. This is interesting as we observe there isn’t an important variation in the probability of leaving across the career, and the hypothesis of an overall geometric distribution for leaving and coming back is quite realistic. We interpret the change of behavior 30 years after admission as artefactual, as almost no one in the population has had the time to arrive at this stage, which can also be seen with the widening of uncertainty.
@@
We now turn to our main quantities of interest, cohort-related coefficients, which are presented in the left pane of Figure 4. The view is quite contrasted between $\gamma^{0}$ and $\gamma^{1}$. On the one hand, it isn’t possible to conclude that the probability of leaving for the private sector $\gamma^{0}$ has significantly changed across the period, as the credible interval well includes zero ($\text{med}=-0.226,I_{90}=[-0.625,0.176]$). However, it is at least possible to state strongly that this coefficient has not increased across time, and has quite possibly decreased. On the other hand, we can state with fair confidence that $\gamma^{1}$ has increased over time, with $\text{med}=1.16$, ($I_{90}=[0.301,1.98]$). We thus conclude that on the one hand, the probability of leaving the public sector has not increased in the period, and may even have decreased; and that the probability of coming back has increased in the period, contrary to the conclusions of rouban_norme_2014.
@@
Dependency with the electoral process. The posterior for the election time effects are reported in the right pane of Figure 4. We observe that the overall size of potential effects is far smaller than for period effects, and so should be interpreted cautiously. On the one hand, we conclude that elections have no impact on $\gamma^{0}$, as the credible interval is very well centered on zero ($\text{med}=-0.0212,I_{90}=[-0.249,0.201]$). On the other hand, however, despite important uncertainty, there seems to be a small positive effect of election time on $\gamma^{1}$ ($\text{med}=0.302,I_{90}=[0.000,0.302]$). In summary, electoral shifts probably do not influence the probability of high civil servants to leave for the private sector, but it might lead to a higher probability of coming back for those who had left.
5.3 Public-private paths across the state
We now move to an even larger population, and consider all top French civil servants ($N=5479$ individuals). Our method and data allow us to describe the heterogeneity of public-private paths across the State and how they interact with gender.
@@ 5.3 Public-private paths across the state
Many social scientists have early insisted on the usefulness of adopting a comparative point of view when describing civil servants populations and public-private paths (charle_pantouflage_1987; kolopp2021pantoufler). We attempt to build upon their results and propose a complete and systematic comparison across the different organizations we mentioned. This has value in the sense that it can qualify statements regarding the most prestigious organizations that are represented in these studies: are they really that special compared to other powerful services? For example, are the Finance Inspection and Treasury that much more affected by revolving doors compared to other services in the Ministry of Economics and Finance, or to other inspections? We can also test statements linked to individual properties: do men have a higher probability to leave for the private sector, and then come back?
@@ 5.3.1 Models
We encode organizational membership using a set of $D$ dichotomous variables $M_{t,1}^{i},...,M_{t,D}^{i}$ corresponding to our $D=11$ selected organizations, that we will include in the covariate matrices $W_{t}^{i}$. Contrary to the previous section, we need to account for the fact that individuals can be affiliated to multiple organizations during their professional career. To remain consistent with our choice to focus on the main organizational identity of individuals, we impose that someone, at any given time, may only be linked to one organization in our set of $D$ organizations. This means we impose $\sum_{d=1}^{D}M_{t,d}^{i}=1$. If an individual is observed to become part of organization $d_{1}$ at time $t_{1}$ and then of organization $d_{2}$ at time $t_{2}$, we will have $M_{t,{d_{1}}}=1$ for $t_{1}\leq t<t_{2}$, and $M_{t,d_{2}}=1$ for $t\geq t_{2}$. This is a simplification, which has the advantage of enabling an analysis that is closest to what we could obtain by performing completely separate analyses for the different public organizations, even if it could be altered to study how having been the member of a given organization can affect behavior once being a member of another one included in the study.
We implement two configurations. Model 1 is a Markov switching model for which we only include an intercept for each group (no general intercept, and no reference organization), in addition to the autoregressive covariates for traces. The autoregressive covariates are interacted with the dummy of every organization. This is our reference model. Model 2 adds to the reference model a dichotomous variable encoding gender ($G^{i}=1\$if $i$ is a woman), for both $\gamma^{0},\gamma^{1}$ and $\lambda$.
@@ 5.3.2 Results
Reconstruction loss. The reconstruction loss obtained by the two models in this study is of $0.279$ and $0.265$, which is higher than for ENA graduates. This can be explained because the inference task is more complex given the higher heterogeneity of the population, and is in line with what could be expected given the nature of the application.
@@
Comparison across organizations. We report the 5%, 50%, and 95% quantiles of the posterior for each of these models. Median estimates of coefficients for Model 1 are reported in Figure 5, alongside their 90% credible interval, expressed in raw probability. Bear in mind that these are average coefficients that do not correspond to local probabilities individuals may have to leave at any given point in time, which are modulated by the observed traces and the local $\lambda$’s. As coefficients for dgfip stand out from the rest in both estimate and uncertainty, we do not report them in the Figure but do so in the text below.
@@
Figure 5: Median and 90% credible interval of coefficients $\gamma^{0}_{\text{group}}$, $\gamma^{1}_{\text{group}}$ for individuals belonging to groups defined in subsection 2.1. Coefficients may be interpreted as average exit ($\gamma^{0}$) or return ($\gamma^{1}$) probabilities, yet it does not mean the model predicts transitions to happen with constant probability.
@@
We first rank organizations depending on the probability that members leave for the private sector. The organizations with the highest departure probabilities are the General Inspection of Finances ($\gamma^{0}_{\text{igf}}=0.031$) and Prefects group ($\gamma^{0}_{\text{cprefet}}=0.029$). In an intermediary situation, we find the Diplomats group ($\gamma^{0}_{\text{cdiplo}}=0.022$), and the General Inspection of Administration ($\gamma^{0}_{\text{iga}}=0.020$) At the lower hand of the spectrum, we find the statistical office ($\gamma^{0}_{\text{insee}}=0.016$), the Cour des Comptes ($\gamma^{0}_{\text{ccomptes}}=0.015$), Treasury ($\gamma^{0}_{\text{dgtresor}}=0.014$), the General Inspection for Social Affairs ($\gamma^{0}_{\text{igas}}=0.013$), and the Conseil d’État ($\gamma^{0}_{\text{ce}}=0.011$). Not shown in the table, the General Direction for Public Finances has the lowest probability ($\gamma^{0}_{\text{dgfip}}=0.006,I_{90}=[0.004,0.009]$). This last very low coefficient means that there are almost no departures from this direction, and when that is the case, civil servants move first to another service included in this study. The overall ranking should be intuitive to the reader accustomed to the question – the model rightly identifies that individuals in the Treasury have a higher probability of leaving for the private sector than individuals in the Inspection for Social Affairs. A seemingly counterintuitive result is the position of igf and ccomptes: these two groups have a similar function (financial auditing of the state), but have very different positions in Figure 5; this observation is in fact aligned with an effect identified in the survey from charle_pantouflage_1987.
@@
We observe interesting contrasts in the probabilities that individuals come back working in public administration after being in the private sector. Inspection of Finances again stands out, with ($\gamma^{1}_{\text{igf}}=0.024$). Next, the Conseil d’État ($\gamma^{1}_{\text{ce}}=0.020$), Inspection of social affairs ($\gamma^{1}_{\text{igas}}=0.019$), Treasury ($\gamma^{1}_{\text{dgtresor}}=0.018$), and Cour des comptes ($\gamma^{1}_{\text{dgtresor}}=0.016$) appear in an intermediate situation. At the lower end of the spectrum, we observe the statistical office ($\gamma^{1}_{\text{insee}}=0.006$), inspection of administration ($\gamma^{1}_{\text{iga}}=0.004$), prefects ($\gamma^{1}_{\text{cprefets}}=0.003$), and diplomats ($\gamma^{1}_{\text{diplo}}=0.002$).
@@
We thus conclude that the two rankings are not convergent and would even be negatively correlated if we excluded igf. However, we find it more accurate to describe this as forming three groups. First, igf constitutes a group on its own, standing out from the rest as having both high probabilities for leaving and coming back. Classical results claiming that moving into the private sector is part of the career advancement process hold well, and we further prove that this is quite distinctive to this organization. We identify a second group of organizations where the probability of leaving is relatively high, but the probability of coming back is rather low: this concerns the statistical office, the inspection of administration, diplomats and préfets. The situation is unsurprising for the first two, which are known to be closer to the private sector than other organizations. For préfets and diplomats, this might be because we study individuals who are simultaneously at the top of certain services but without as high position security as in other organizations (such as Conseil d’État, where people can keep their status): people (involuntarily) leaving could find themselves unable to come back. Finally, we find a group of organizations where the probability to leave is comparatively low and the probability of coming back high: Conseil d’État, Inspection of Social Affairs, Cour des Comptes. One could also place Treasury in this group, but parameter uncertainty (especially on $\gamma^{1}$) requires caution. This concerns organizations for which members acquire life-long status that can help them come back after leaving; and also for which clear career advancement schemes exist in the public sector.
Usefulness of combining the two data sources We repeated the analysis using only the LinkedIn data, and show in Figure 6 the posterior distribution of $\gamma^{0}$ per organization. Notice that for many groups, the posterior distribution is very different, which shows that using only the LinkedIn data leads to bias. Similar effects were observed for other parameters of interest. It is not possible to perform inference using only the trace data: the model parameters are not identifiable, and the MCMC algorithm therefore does not converge.
Figure 6: Comparison of results between initial study on various groups (top) and study where only individuals with a LinkedIn profile are included (bottom). Coefficients are shown in average probability scale. The posterior distributions are very different, indicating that the results on the bottom row are biased.
@@
Interaction with gender We now consider the effect of the gender covariate. Figure 6 displays the posterior distribution of the odds ratio of $\gamma^{0}$ for women versus men. A distribution shifted to the right indicates that women have a higher probability of leaving for the private sector than men, and vice-versa when shifted to the left. A distribution centered on 1 indicates no significant gender effect.
@@
Figure 7: Posterior distribution of $\frac{\gamma^{0}_{\text{group},W}/(1-\gamma^{0}_{\text{group},W})}{\gamma^{0}_{\text{group},M}/(1-\gamma^{0}_{\text{group},M})}$, the odds-ratio of the differential of exit probabilities between women and men expressed, for each of the nine organizations defined in subsection 2.1. When the posterior is concentrated on the left-hand side, we can conclude that women in that group have a lower probability than men of exiting the public sector.
@@
Not all organizations have the same profile of differentiated practices between men and women, but a broad conclusion can nonetheless be inferred: women have a lower probability than men to leave for the private sector. For certain organizations, the difference is strong and identifiable: Cour des Comptes, Inspection of Administration, Inspection of Social Affairs (odds ratios below 0.7). The difference is smaller but still observable for Inspection of Finances, Préfets, and Diplomats. The situation is undecided for Conseil État and the Statistical Office, the latter suggesting a small effect of women having a lower probability of leaving. Treasury here stands out in terms of uncertainty (as there are fewer than 200 individuals to work on and a small proportion of women), but the evidence tends to point at a higher probability of leaving for women.
6 Discussion